# OE_^ OmOE_^ OmOE_^ Omm!OE_^ mmmmOE_^

### v1

@@ 1 Introduction
Multimodal Large Language Models (MLLMs) bai2023qwenvl ; wang2024qwen2vl ; bai2025qwen25vl ; chen2024internvl ; chen2024internvl1.5 ; team2024internvl2 ; chen2024expanding2.5 ; yao2024minicpm ; agrawal2024pixtral ; Emu2 ; wang2024emu3 ; chen2025janus have demonstrated remarkable progress in recent years, excelling at tasks such as visual question answering fu2024mme ; yu2023mm ; yue2024mmmu , text-to-image generation ghosh2023geneval ; hu2024ella , and video understanding li2024mvbench ; zhou2024mlvu ; wang2024lvbench ; fu2024video ; chen2024we . Despite these advances, existing MLLMs still struggle with spatial reasoning mayer2025ivispar ; tang2025lego —a fundamental aspect of human intelligence. Given this critical gap in emulating human-like spatial cognition, it becomes imperative to develop a targeted framework for assessing such aspect. To this end, this paper takes the challenging task of Transformation Driven Visual Reasoning (TVR) hong2021transformation as a starting point to investigate MLLMs’ capabilities in spatial understanding. By comparing two given images containing multiple objects, TVR involves a series of continuous thinking and reasoning steps to derive the correct answer, including identifying which object is transformed and what transformation is applied to this object. Moreover, objects in the two images may be presented in different viewpoints, which further complicates the task and poses difficulties for existing MLLMs. Our experiments show that, the powerful closed-source commercial model GPT-4o hurst2024gpt achieves only 23.5% accuracy on this task, further demonstrating that this problem poses a significant challenge to existing MLLMs.
@@
To address these challenges, we propose to integrate dense reward plus punishment into GRPO, thereby constructing a new method dubbed Spacial TrAnsformation Reasoning with R1 (STAR-R1). STAR-R1 is built upon two core designs. On the one hand, it introduces dense rewards by accounting for partial answer correctness, progressively assigning fine-grained rewards to three-level answers. Specifically, for the first-level answer, the model identifies the objects with modified attributes, receiving a minimal reward. For the second-level answer, it correctly predicts the changed attributes, earning a medium reward. For the third-level answer, the model accurately identifies the full object-attribute-transformation triplet, receiving the highest reward. On the other hand, STAR-R1 penalizes the model for either excessively enumerating object-attribute-change triplets to maximize rewards or remaining inactive to avoid incorrect predictions, ensuring active exploration without shortcuts. This mechanism design instills into the model the learning philosophy that prioritizes attempting potentially incorrect answers over overlooking possible correct ones.
@@ 2 Related Work
Multimodal Reasoning Model. Research on the reasoning capabilities of MLLMs has attracted significant attention in recent years. Early studies primarily employed Chain-of-Thought (CoT) fine-tuning to enhance model reasoning abilities (xu2024llava, ; xu2025redstar, ), while the recent success of DeepSeek-R1 (guo2025deepseek, ) has promoted the application of rule-based Reinforcement Learning in this field (chen2025r1v, ; shen2025vlm, ). Notably, MM-Eureka (meng2025mm, ) adopts an online filtering strategy and a two-stage multimodal RL training framework to achieve significant improvements in mathematical reasoning tasks; LMM-R1 (peng2025lmm, ), which enhances multimodal reasoning capabilities through a two-phase training paradigm combining text-only RL and multimodal RL; and Video-R1 (feng2025video, ), which achieves breakthroughs in video reasoning tasks via a two-stage approach involving supervised fine-tuning (SFT) followed by RL training. However, existing methods exhibit several limitations: (1) most rely on naive GRPO strategies, (2) they predominantly adopt two-stage training paradigms, and (3) there is a lack of task-specific customization and in-depth exploration of pure RL training (e.g., like DeepSeek-R1-Zero). To address these gaps, we introduces a tailored rule-based RL strategy for the TVR task and proposes a single-stage, pure RL training paradigm. Our approach not only significantly enhances multimodal reasoning performance but also provides novel theoretical insights into RL training mechanisms.

### v3

@@ 1 Introduction
Multimodal Large Language Models (MLLMs) bai2023qwenvl ; wang2024qwen2vl ; bai2025qwen25vl ; chen2024internvl ; chen2024internvl1.5 ; team2024internvl2 ; chen2024expanding2.5 ; yao2024minicpm ; agrawal2024pixtral ; Emu2 ; wang2024emu3 ; chen2025janus have demonstrated remarkable progress in recent years, excelling at tasks such as visual question answering fu2024mme ; yu2023mm ; yue2024mmmu , text-to-image generation ghosh2023geneval ; hu2024ella , and video understanding li2024mvbench ; zhou2024mlvu ; wang2024lvbench ; fu2024video ; chen2024we . Despite these advances, existing MLLMs still struggle with spatial reasoning mayer2025ivispar ; tang2025lego ; jia2025omnispatial ; liu2025ir3d —a fundamental aspect of human intelligence. Given this critical gap in emulating human-like spatial cognition, it becomes imperative to develop a targeted framework for assessing such aspect. To this end, this paper takes the challenging task of Transformation Driven Visual Reasoning (TVR) hong2021transformation as a starting point to investigate MLLMs’ capabilities in spatial understanding. By comparing two given images containing multiple objects, TVR involves a series of continuous thinking and reasoning steps to derive the correct answer, including identifying which object is transformed and what transformation is applied to this object. Moreover, objects in the two images may be presented in different viewpoints, which further complicates the task and poses difficulties for existing MLLMs. Our experiments show that, the powerful closed-source commercial model GPT-4o hurst2024gpt achieves only 23.5% accuracy on this task, further demonstrating that this problem poses a significant challenge to existing MLLMs.
@@
To address these challenges, we propose to integrate dense reward plus punishment into GRPO, thereby constructing a new method dubbed Spatial TrAnsformation Reasoning with R1 (STAR-R1). STAR-R1 is built upon two core designs. On the one hand, it introduces dense rewards by accounting for partial answer correctness, progressively assigning fine-grained rewards to three-level answers. Specifically, for the first-level answer, the model identifies the objects with modified attributes, receiving a minimal reward. For the second-level answer, it correctly predicts the changed attributes, earning a medium reward. For the third-level answer, the model accurately identifies the full object-attribute-transformation triplet, receiving the highest reward. On the other hand, STAR-R1 penalizes the model for either excessively enumerating object-attribute-change triplets to maximize rewards or remaining inactive to avoid incorrect predictions, ensuring active exploration without shortcuts. This mechanism design instills into the model the learning philosophy that prioritizes attempting potentially incorrect answers over overlooking possible correct ones.
@@ 2 Related Work
Multimodal Reasoning Model. Research on the reasoning capabilities of MLLMs has attracted significant attention in recent years. Early studies primarily employed Chain-of-Thought (CoT) fine-tuning to enhance model reasoning abilities (xu2024llava, ; xu2025redstar, ), while the recent success of DeepSeek-R1 (guo2025deepseek, ) has promoted the application of rule-based Reinforcement Learning in this field (chen2025r1v, ; shen2025vlm, ; yang2025deepcritic, ; fan2025sophiavl, ; liang2025swsselfawareweaknessdrivenproblem, ; li2025temporal, ; wang2025solidgeomeasuringmultimodalspatial, ; yang2025towards, ; song2025maniplvmr1reinforcementlearningreasoning, ). Notably, MM-Eureka (meng2025mm, ) adopts an online filtering strategy and a two-stage multimodal RL training framework to achieve significant improvements in mathematical reasoning tasks; LMM-R1 (peng2025lmm, ), which enhances multimodal reasoning capabilities through a two-phase training paradigm combining text-only RL and multimodal RL; and Video-R1 (feng2025video, ), which achieves breakthroughs in video reasoning tasks via a two-stage approach involving supervised fine-tuning (SFT) followed by RL training. However, existing methods exhibit several limitations: (1) most rely on naive GRPO strategies, (2) they predominantly adopt two-stage training paradigms, and (3) there is a lack of task-specific customization and in-depth exploration of pure RL training (e.g., like DeepSeek-R1-Zero). To address these gaps, we introduces a tailored rule-based RL strategy for the TVR task and proposes a single-stage, pure RL training paradigm. Our approach not only significantly enhances multimodal reasoning performance but also provides novel theoretical insights into RL training mechanisms.
3 Our Method: STAR-R1
@@ 5 Conclusion
Appendix A Dataset and Experimental Setup
We first clean the TRANCE dataset [22] by removing samples containing redundant transformations. Then, we randomly sample 9,000 and 4,500 samples as the training set and test set, respectively, ensuring that the number of samples corresponding to each transformation length (ranging from 1 to 4) is equally distributed in both the training and test sets.
We employ Qwen2.5-VL-7B as our base model and utilize vLLM [26] as the training framework. All experiments (including RL and SFT) are conducted on a single node equipped with 8 $\times$ H20 GPUs. Except for the training strategy mentioned in Appendix C, all other training runs adopt a single-stage procedure with 2 epochs, a batch size of 1 per GPU, and gradient accumulation over 2 steps, yielding a total of 1,286 training steps.
Appendix B Training Curves
Figure 5: Training curves about accuracy reward and format reward.
We illustrate the evolution of Total Accuracy for STAR-RL and STAR-SFT models on OOD test data across training steps. The training dynamics plotted in  Fig. 7 further show that SFT quickly plateaus around 1,200 steps on OOD data and eventually degrades with further training, whereas RL maintains steady performance growth throughout the training process, suggesting its unique ability to progressively unlock the model’s reasoning capabilities for reasoning in complex scenarios.
Additionally, we plot further curves to analyze the training dynamics of STAR-R1. As shown in the right panel of  Fig. 5, we present the format reward curve of the STAR-R1. It can be observed that for models with strong foundational capabilities, such as Qwen-2.5-VL-7B, they already exhibit excellent instruction-following performance in the TVR task from the beginning, with most responses adhering to the required format. Thus, no additional customization for format reward is necessary. Moreover, in the left panel of  Fig. 5, we illustrate the accuracy reward curve. The model’s accuracy reward steadily increases with training progression, ultimately stabilizing at a high value. This indicates that our designed reward function effectively encourages the model to engage in accurate, proactive, and efficient exploration, progressively unlocking the multimodal reasoning capabilities of the base model.
Figure 6: Dynamics of the total accuracy as well as the attribute accuracy for the four types throughout the training process.
As shown in  Fig. 6, these five curves illustrate the dynamics of the Total Accuracy as well as the Attribute Accuracy for the four types throughout the training process. All accuracy curves exhibit a steady increase during training and eventually stabilize at a high level of prediction accuracy. This demonstrates that our STAR-R1 effectively guides the model to efficiently explore the solution space and gradually develop stronger reasoning capabilities.
Figure 7: Accuracy curves of RL and SFT.
Table 6: Comparison of RL vs. SFT performance on ID and OOD settings.
[…]
Method TAcc Diff NDiff TAcc Diff NDiff
 STAR-SFT 84.2 0.22 0.08 30.9 1.65 0.83
STAR-SFT&RL 87.7 0.19 0.07 36.5 1.53 0.71
STAR-R1 76.3 0.38 0.14 53.9 0.96 0.39
[…]
Appendix C Training STAR-SFT with RL
We additionally apply RL training to the SFT model. In Row 2 of  Table 6, results demonstrate that further training the model with RL not only achieves continuous improvement in ID performance but, more importantly, exhibits significant progress on OOD tasks. This indicates that RL successfully breaks through the performance bottleneck encountered during SFT training on OOD data. Furthermore, we observed that the two-phase training approach of STAR-SFT&RL do not ultimately achieve the same performance on OOD data as pure RL training. We hypothesize that this may be due to imitation learning in the SFT phase locking the model’s reasoning patterns, such that the subsequent RL training phase using the same amount of data can only marginally adjust the model’s reasoning. This adjustment leads the model to attempt comparing more objects.
To validate this conjecture, we further analyze the difference between the number of objects compared in each response and the total number of objects in the scene. The average difference for STAR-SFT is -3.3, while for STAR-SFT&RL, it is -3.0. This indicates a slight increase in the number of objects compared during the reasoning process of the STAR-SFT&RL. Consequently, this results in a modest improvement in OOD accuracy, though it still falls short of the performance achieved by the pure RL training.
Appendix D More Cases
In this section, we provide a detailed presentation of our problem prompt (shown in Fig. 8) along with additional case studies. The {ObjectFeature} in Fig. 8 will be substituted by all features corresponding to each object in the initial image of every sample. An example is provided below: {idx: 0; color: yellow; material: metal; shape: cylinder; size: medium}, {idx: 1; color: gray; material: rubber; shape: sphere; size: medium}, {idx: 2; color: blue; material: rubber; shape: cylinder; size: medium}, {idx: 3; color: yellow; material: metal; shape: sphere; size: medium}, {idx: 4; color: blue; material: rubber; shape: cube; size: large}, {idx: 5; color: brown; material: rubber; shape: sphere; size: medium}, {idx: 6; color: brown; material: rubber; shape: sphere; size: large}.
The case studies shown in Figs. 9, 10 and 11 further substantiate the argument presented in Section 4.2 of the main paper, namely that STAR-SFT merely engages in imitation learning and does not perceive changes in point of view. It erroneously associates objects located at the same positions in the initial and final images (referring to their positions within the images, not their positions in the real-world scene) as the same object, thereby significantly impairing the model’s reasoning performance.
In contrast, STAR-R1 demonstrates anthropomorphic behavior by systematically comparing the states of all objects between the initial and final images during its reasoning process to ascertain their correspondence. Consequently, without any explicit mention of viewpoint changes in the problem prompts, STAR-R1 successfully detects the perspective shift and achieves a significantly higher answer accuracy than STAR-SFT.
Since the STAR-SFT model fails to perceive changes in viewpoint, it mistakenly assumes that objects at the same position in two images are the same object. In  Fig. 9, the STAR-SFT model incorrectly identifies objects numbered 0, 3, and 5 as objects numbered 3, 5, and 6. In  Fig. 10, it misidentifies objects numbered 1, 0, and 5 as objects numbered 0, 5, and 4. In  Fig. 11, the model confuses objects numbered 3, 4, and 0 with objects numbered 7, 3, and 4. In contrast, by comparing the states of all objects in both images, STAR-R1 correctly matches the same objects across the two images and ultimately outputs the correct answer.
Figure 8: The problem prompt.
Figure 9: A case study comparing the reasoning processes of STAR-SFT and STAR-R1. Blue indicates correct answers while red denotes incorrect ones.
Figure 10: A case study comparing the reasoning processes of STAR-SFT and STAR-R1. Blue indicates correct answers while red denotes incorrect ones.
Figure 11: A case study comparing the reasoning processes of STAR-SFT and STAR-R1. Blue indicates correct answers while red denotes incorrect ones.
Appendix E Limitations
This work studies the problem of concurrent object transformations at a single timestamp. In contrast, real-world scenarios often involve temporally extended sequences of transformations with latent dependencies. Extending our approach to such long-term settings, accounting for temporal causality and state memory, is a key direction for future research.